OpenStudy (anonymous):

what is the LCD for 50 and 30?

OpenStudy (amistre64):

the smallest common multiple .... can be determined by breaking them downinto their prime products; then seeing what you have to include to equal them out

OpenStudy (amistre64):

the prime factorizations are: 50 = 2*5*5 30 = 2*3*5 one manner to determine how to fill in the rest is to line them up so that like values are above like values 2 5 5 2 3 5 now drop everything down -------- 2 3 5 5 these are the prime factors of the LCD

OpenStudy (anonymous):

so 150 is the answer. thank you :)

OpenStudy (amistre64):

i like 150, good luck ;)

OpenStudy (anonymous):

thanks :)

OpenStudy (unklerhaukus):

so we take the product of the disjunct of the factors of each number

OpenStudy (amistre64):

i thought about trying to state it in set terms as well, but most proper sets dont accept multiples of the same element.

OpenStudy (amistre64):

if we consider the factors to be distinct, regardless of multiplicity, then yes

OpenStudy (unklerhaukus):

hmmm